# Teaching Robots Like Dogs: Learning Agile Navigation from Luring, Gesture, and Speech

### v1

@@ III Problem Formulation
As shown in Figure 2, we decompose the problem into two subsystems: a high-level navigation module and a low-level locomotion controller. The navigation module, denoted as $\pi_{g}$, serves as a high-level System 2 planner that interprets human intent and predicts navigation goals $\mathbf{g}$, represented as $\mathbf{g}=\pi_{g}(\mathbf{c},\mathbf{x},\bm{\rho})$. The locomotion controller, represented as $\mathbf{u}=\pi_{u}(\mathbf{g},\mathbf{x},\bm{\rho})$, functions as a fast and reactive System 1 module that executes agile behaviors to overcome obstacles. By integrating these two modules, robot is controlled directly by high-level interaction commands.
@@ IV-A Locomotion controller
As a preliminary stage, we develop a locomotion controller for low-level agility of the robot. As shown in the Stage 0 from Figure 3, the locomotion controller consists of a velocity planner and a velocity tracker. The velocity planner converts the navigation goal $\mathbf{g}$ into a velocity command guiding the robot toward its target. In more detail, it outputs a heading velocity command $v_{\text{com}}=K\|\mathbf{g}^{xy}-\mathbf{p}^{xy}\|_{2}$ and an rotation command $\phi_{\text{com}}=\text{yaw}(R^{\top}(\mathbf{g}-\mathbf{p}))$, where $\mathbf{p}\in\mathbb{R}^{3}$ is the robot’s base position, $R\in\mathrm{SO}(3)$ is the robot’s rotation matrix and $(\cdot)^{xy}$ indicates ground-projected coordinates.
@@ IV-D Progressive goal cueing
It is crucial to provide interaction commands and navigation goals that align with the robot’s current state during training, as misalignment can lead to inconsistent supervision. One naive approach is to feed gesture and verbal commands to the policy according to their original timestamps. However, during early training, the policy often progresses more slowly through tasks, leading to new commands being issued before the robot has completed the preceding behavior. Since the interaction data were collected by issuing new commands in response to the robot’s task progress, this navie replay strategy can introduce state–command misalignment.
@@ V-B4 LURE (Ours)
Succuess Rate $\uparrow$ Navigation Error $\downarrow$ Succuess Rate $\uparrow$ Navigation Error $\downarrow$ Succuess Rate $\uparrow$ Navigation Error $\downarrow$
@@ V-D Adaptation to novel user
We evaluate whether our system can generalize to novel users through quick adaptation. This is essential because different users often employ different interaction cues to command the robot for the same task. We first train a model using the initial interaction dataset from a single user described in Section V-A, and then fine-tune it using an average of 4.5 minutes of data collected from each new subject. For evaluation, measure the success rate and navigation error in the reconstructed simulation environment, following the same protocol described in Section V-C.
@@ VI Limitations
Finally, our current vision system is restricted to controlled environments, as the objects are modeled as circular or square primitives with pre-measured dimensions, and the human wears a motion capture suit for pose estimation. Camera-based human pose estimation and scene reconstruction [16, mihajlovic2025volumetricsmpl] could potentially remove these assumptions and enable operation in unstructured, real-world scenarios.
@@ VII Conclusion and Future Work
We presented a framework that enables robots to learn and execute navigation skills in real time through natural human interaction. By reconstructing interaction scenes and augmenting data within these environments, our approach enables policies to robustly recover from distributional shifts across diverse interaction conditions. Experimental results demonstrate that the proposed framework effectively integrates verbal and gestural cues, enabling robust control across a range of tasks and environments. Moreover, our findings highlight the complementary nature of speech and gesture, showing that their combination provides a powerful interface for teaching and controlling robots to execute context-appropriate behaviors, including agile maneuvers such as jumping over obstacles. Overall, this work advances an intuitive and practical system that directly maps high-level human intent to executable robot actions. In future work, we will explore extending our framework beyond navigation to more general whole-body behaviors, such as locomotion manipulation, by grounding multimodal commands using a similar data augmentation strategy.

### v2

@@ III Problem Formulation
As shown in Figure 2, we decompose the problem into two subsystems: a high-level navigation module and a low-level locomotion controller. The navigation module, denoted as $\pi_{g}$, serves as a high-level System 2 planner that interprets human intent and predicts navigation goals $\mathbf{g}$, represented as $\mathbf{g}=\pi_{g}(\mathbf{c},\mathbf{x},\bm{\rho})$. The locomotion controller, represented as $\mathbf{u}=\pi_{u}(\mathbf{g},\mathbf{x},\bm{\rho})$, functions as a fast and reactive System 1 module that executes agile behaviors to overcome obstacles. By integrating these two modules, the robot is controlled directly by high-level interaction commands.
@@ IV-A Locomotion controller
As a preliminary stage, we develop a locomotion controller for low-level agility of the robot. As shown in the Stage 0 from Figure 3, the locomotion controller consists of a velocity planner and a velocity tracker. The velocity planner converts the navigation goal $\mathbf{g}$ into a velocity command guiding the robot toward its target. In more detail, it outputs a heading velocity command $v_{\text{com}}=K\|\mathbf{g}^{xy}-\mathbf{p}^{xy}\|_{2}$ and a rotation command $\phi_{\text{com}}=\text{yaw}(R^{\top}(\mathbf{g}-\mathbf{p}))$, where $\mathbf{p}\in\mathbb{R}^{3}$ is the robot’s base position, $R\in\mathrm{SO}(3)$ is the robot’s rotation matrix and $(\cdot)^{xy}$ indicates ground-projected coordinates.
@@ IV-D Progressive goal cueing
It is crucial to provide interaction commands and navigation goals that align with the robot’s current state during training, as misalignment can lead to inconsistent supervision. One naive approach is to feed gesture and verbal commands to the policy according to their original timestamps. However, during early training, the policy often progresses more slowly through tasks, leading to new commands being issued before the robot has completed the preceding behavior. Since the interaction data were collected by issuing new commands in response to the robot’s task progress, this naive replay strategy can introduce state–command misalignment.
@@ V-B4 LURE (Ours)
Success Rate $\uparrow$ Navigation Error $\downarrow$ Success Rate $\uparrow$ Navigation Error $\downarrow$ Success Rate $\uparrow$ Navigation Error $\downarrow$
@@ V-D Adaptation to novel user
We evaluate whether our system can generalize to novel users through quick adaptation. This is essential because different users often employ different interaction cues to command the robot for the same task. We first train a model using the initial interaction dataset from a single user described in Section V-A, and then fine-tune it using an average of 4.5 minutes of data collected from each new subject. For evaluation, we measure the success rate and navigation error in the reconstructed simulation environment, following the same protocol described in Section V-C.
@@ VI Limitations
Finally, our current vision system is restricted to controlled environments, as the objects are modeled as circular or square primitives with pre-measured dimensions, and the human wears a motion capture suit for pose estimation. Camera-based human pose estimation and scene reconstruction [17, 12] could potentially remove these assumptions and enable operation in the wild.
@@ VII Conclusion and Future Work
We presented a framework that enables robots to learn and execute navigation skills in real time through natural human interaction. By reconstructing interaction scenes and augmenting data within these environments, our approach enables policies to robustly recover from distributional shifts across diverse interaction conditions. Experimental results demonstrate that the proposed framework effectively integrates verbal and gestural cues, enabling robust control across a range of tasks and environments. Moreover, our findings highlight the complementary nature of speech and gesture, showing that their combination provides a powerful interface for teaching and controlling robots to execute context-appropriate behaviors, including agile maneuvers such as jumping over obstacles. Overall, this work advances an intuitive and practical system that directly maps high-level human intent to executable robot actions. In future work, we will explore extending our framework beyond navigation to more general whole-body behaviors, such as loco-manipulation, by grounding multimodal commands using a similar data augmentation strategy.